OpenStudy (anonymous):

Determine solution set of the inequality \[|x-1|\geq 3|x+1|\]

OpenStudy (anonymous):

i think you need to work in cases if \(x>1\) then \(|x-1|=x-1\) and \(|x+1|=x+1\) first job it to solve \[x-1\geq 3(x+1)\]

OpenStudy (anonymous):

if \(x<-1\) then \(|x-1=1-x\) and \(|x+1|=-x-1\) etc

OpenStudy (anonymous):

the solution is -2<x<-1/2 but i don't know how we found it..

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

if \(x>1\) then both \(x-1\) and \(x+1\) are positive, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which means they are equal to their absolute values in other words \(|x+1|=x+1\) and \(|x-1|=x-1\)

OpenStudy (anonymous):

so if \(x>1\) you need to solve \[x-1\geq 3(x+1)\]

OpenStudy (anonymous):

we solve in a couple of steps \[x-1\geq 3x+3\]\[-1\geq 2x+3 \\-4\geq 2x \\x<-2\] but here we have a problem because we were assuming that \(x>1\) but came up with \(x<-2\) so if \(x>1\) there is NO solution

OpenStudy (anonymous):

now we try again with \[x<-1\] but this time both \(x-1\) and \(x+1\) are negative, and so \[|x+1|=-x-1\] and \[|x-1|=-x+1\]

OpenStudy (anonymous):

we solve \[-x+1\geq 3(-x-1)\] in the same number of steps \[-x+1\geq -3x-3\] \[2x+1\geq -3\] \[2x\geq -4\] \[x\geq -2\]

OpenStudy (anonymous):

that means if \(x<-1\) then also \(x>-2\) so we know we are good from \(-2\) up to \(-1\)

OpenStudy (anonymous):

thanks Satellite, what about 1/2 we still didn't get it..

OpenStudy (anonymous):

(sorry i got internet problem before)